# Responsive Noise-Relaying Diffusion Policy: Responsive and Efficient Visuomotor Control

### v1

@@ 6.5 Ablation Study
Noise Scheduling Scheme We conduct a comprehensive study to evaluate how different noise scheduling schemes impact the performance of the Responsive Noise-Relaying Diffusion Policy. As shown in Table 5, the results demonstrate that relying solely on either a linear schedule or a random schedule significantly reduces task success rates. This highlights the importance of integrating mixture scheduling to fully exploit the potential of the noise-relaying buffer design. By combining the two schedules, our model effectively utilizes different noise levels, enhancing the robustness and adaptability of action generation.
Table 5: Ablation study on noise scheduling scheme during training (Section 5.2). Numbers represent average success rates ($\uparrow$). Numbers in parenthesis indicate the performance drop after removing key component of our Responsive Noise-Relaying Diffusion Policy.
@@
Noise-Relaying Buffer Initialization We conduct an ablation study to evaluate the effectiveness of our initialization scheme. Pure noise directly uses a fully noisy buffer as the initialization for subsequent operations. As shown in Table 6, laddering initialization achieves significant performance gains compared to pure noise. This result highlights the critical role of our initialization scheme in enabling strong and robust performance.
Table 6: Ablation study on noise-relaying buffer initialization (Section 5.2).
@@
Noise-Relaying Buffer Capacity We also evaluate the impact of the noise-relaying buffer capacity to the performance, the only tuned hyperparameter in our approach. As shown in Figure 5, on the Adroit Relocate task, RNR-DP achieves peak performance of $58.5\%$ at a buffer capacity of 84 and maintains strong performance within the range of 56 to 92, demonstrating a wide tolerance. Performance declines when the buffer capacity is too small (e.g., 48) or too large (e.g., 100). Notably, Diffusion Policy achieves only $42.2\%$, while RNR-DP with buffer capacities between 48 and 100 consistently outperforms it. This highlights that for a single task, the noise-relaying buffer offers robust performance over a broad range and is not difficult to tune. Empirically, starting with a value between 56 and 84 and making adjustments suffices most cases.
Model Prediction Type In addition, we investigate the impact of the model’s prediction type. As shown in Table 7, models predicting the added noise component outperform those directly predicting the action sequence. This finding aligns with the common practice of using noise prediction in diffusion models within the vision domain (Ho et al., 2020).
Table 7: Ablation study on model prediction type (Section 5.2).
@@ 7 Conclusion and Future Work
Limitations. A limitation of this work is the lack of additional experiments to demonstrate whether our method preserves behavioral diversity compared to Diffusion Policy. This remains an area for future research.
@@ A.1 Task Descriptions
We consider a total of 9 continuous control tasks from 3 benchmarks: ManiSkill2 (Gu et al., 2023), ManiSkill3 (Tao et al., 2024), and Adroit (Rajeswaran et al., 2017). This section provides detailed task descriptions on overall information, task difficulty, object sets, state space, and action space. Some task details are listed in Table 8.
@@ A.1.1 ManiSkill2 Tasks
  • Overall Description: Pick up a red cube and place it onto a green one. See Figure 6 for episode visualization.
@@
  • Task Difficulty: This task needs to handle object variations. See Figure 7 for episode visualization.
@@
  • Overall Description: A dual-arm mobile robot needs to push a swivel chair to a target location on the ground (indicated by a red hemisphere) and prevent it from falling over. The friction and damping parameters for the chair joints are randomized. See Figure 8 for episode visualization.
@@ A.1.2 ManiSkill3 Tasks
  • Overall Description: It is a simulated version of the real-world push-T task from Diffusion Policy: https://diffusion-policy.cs.columbia.edu/. In this task, the robot needs to precisely push the T-shaped block into the target region, and move the end-effector to the end-zone which terminates the episodes. The success condition is that the T block covers 90% of the 2D goal T’s area. See Figure 9 for episode visualization.
@@
  • Overall Description: A task where the objective is to push and roll a ball to a goal region at the other end of the table. The success condition is that The ball’s xy position is within goal radius (default 0.1) of the target’s xy position by euclidean distance. See Figure 10 for episode visualization.
@@ A.1.3 Adroit Tasks
  • Overall Description: The environment is based on the Adroit manipulation platform, a 28 degree of freedom system which consists of a 24 degrees of freedom ShadowHand and a 4 degree of freedom arm. The task to be completed consists on undoing the latch and swing the door open. See Figure 11 for episode visualization.
@@
  • Overall Description: The environment is based on the Adroit manipulation platform, a 28 degree of freedom system which consists of a 24 degrees of freedom ShadowHand and a 4 degree of freedom arm. The task to be completed consists on repositioning the blue pen to match the orientation of the green target. See Figure 12 for episode visualization.
@@
  • Overall Description: The environment is based on the Adroit manipulation platform, a 28 degree of freedom ShadowHand and a 4 degree of freedom arm. The task to be completed consists on picking up a hammer with and drive a nail into a board. See Figure 13 for episode visualization.
@@
  • Overall Description: The environment is based on the Adroit manipulation platform, a 30 degree of freedom system which consists of a 24 degrees of freedom ShadowHand and a 6 degree of freedom arm. The task to be completed consists on moving the blue ball to the green target. See Figure 14 for episode visualization.
@@ A.2 Demonstrations
This subsection provides the details of demonstrations used in our experiments. See Table 9. ManiSkill2 and ManiSkill3 demonstrations are provided in Gu et al. (2023) and Tao et al. (2024), and Adroit demonstrations are provided in Rajeswaran et al. (2017).
@@ B.1 Noise-Relaying Diffusion Policy Inference
We summarize the inference pseudo-code of our RNR-DP in Algorithm 1.
@@
3:     $\mathbf{Q}_{t}$ $\leftarrow\varepsilon_{\theta}(\mathbf{\tilde{Q}}_{t};\mathbf{O}_{t},\{1,\cdots,f\})$triangleright $\varepsilon_{\theta}$ is trained using $f$ noise levels
4:     $\mathbf{a}_{t}^{(0)}$ $\leftarrow\mathbf{Q}_{t}\mathrm{.pop}(0)$triangleright $\mathbf{a}_{t}^{(0)}$ is a clean action (fully denoised)
5:     $\mathbf{\tilde{Q}}_{t}$ $\leftarrow\mathbf{Q}_{t}\mathrm{.push}(\mathbf{z})$triangleright $\mathbf{z}$ is a random noisy action sampled from $\mathcal{N}(\bm{0},\mathbf{I})$
6:     $\mathbf{O}_{t}$ $\leftarrow\mathrm{env.step}(\mathbf{a}_{t}^{(0)})$triangleright executes $\mathbf{a}_{t}^{(0)}$ and envrionment updates observation
7:     $\hphantom{1}t$ $\leftarrow t+1$triangleright update timestep for the next iteration
@@ B.2 Noise-Relaying Diffusion Policy Training
We summarize the training pseudo-code of our RNR-DP in Algorithm 2.
@@
6:         $\mathbf{k}=\{k_{1}=1,\cdots,k_{f}=f\}$triangleright linear schedule
@@
8:         $\mathbf{k}=\{k_{1}\sim\mathrm{Unif}(\{1,\cdots,f\}),\cdots,k_{f}\sim\mathrm{Unif}(\{1,\cdots,f\})\}$triangleright random schedule
@@
10:         $\hat{\mathbf{a}}_{j}=\sqrt{\bar{\alpha}_{t_{j}}}\mathbf{a}_{j}+\sqrt{1-\bar{\alpha}_{t_{j}}}\bm{\epsilon}_{j}$triangleright perturbe each $\mathbf{a}_{j}$ independently
@@
13:     $\quad\ \ \nabla_{\theta}\|\bm{\epsilon}-\varepsilon_{\theta}(\hat{\mathbf{A}};\color[rgb]{0.48,0.65,0.7}{\mathbf{O},\mathbf{k}}\color[rgb]{0,0,0}{)\|}$ ) ∥ $\triangleright$ noise-aware conditioning
@@ B.3 Policy Architecture
We build our RNR-DP on top of the UNet-based architecture of Diffusion Policy (Chi et al., 2023). The model includes 2 downsampling modules and 2 upsampling modules with each module containing 2 residual blocks. The residual block consists of 1D temporal convolutions (Conv1d), group normalizations (GN), and Mish activation layers. The encoded noise-aware conditioning data (Section 5.2) is fused into each residual block through the FiLM transformation (Perez et al., 2018). The raw conditioning data is of shape $R^{f\times(C_{\mathrm{emb}}+C_{\mathrm{state}})}$ for state policies and of shape $R^{f\times(C_{\mathrm{emb}}+C_{\mathrm{visual}}+C_{\mathrm{state}})}$ for visual policies. See Figure 15 for the visualization of a visual policy. We follow the UNet denoiser design for the observation that transformer-based policies are more sensitive to hyperparameters and often require more tuning (Chi et al., 2023). The choice of policy architecture is orthogonal to our method and we believe our design would also improve this policy class.
@@ B.4.1 Key Hyperparameters of RNR-DP
We summarize the key hyperparameters of RNR-DP in Table 10. The observation horizon $T_{o}$ and noise-relaying buffer capacity $f$ for each task is listed in Table 11. The number of trainable parameters for each task is listed in Table 12.
@@ B.4.2 Key Hyperparameters of Diffusion Policy
We summarize the key hyperparameters of Diffusion Policy in Table 13. The observation horizon $T_{o}$, action executation horizon $T_{a}$ and action prediction horizon $T_{p}$ for each task are listed in Table 14.
@@ C.1 Empirical Comparison with Acceleration Methods on Visual Observations
We summarize the results of vision-based experiments in Table 15. As shown in Table 15, our RNR-DP ourperforms all DDIM variations and CP variations and particularly has an overall improvement over 8-step DDIM by $6.9\%$, over 8-step-chaining CP by $3.4\%$.
Table 15: Evaluation on simpler tasks (Regular Group) not requiring responsive control from ManiSkill and Adroit benchmarks (Visual Observations). We follow our evaluation metric and report values under the same settings as in Table 4. Tasks in which none of the methods achieve a reasonable success rate under visual observations are omitted.
@@ C.2 Comparsion with Streaming Diffusion Policy (SDP)
Streaming Diffusion Policy (SDP) (Høeg et al., 2024) is a recent advancement over Diffusion Policy that stays close to our approach. In this section, we compare our method with SDP in terms of motivation (Section C.2.1), methodology (Section C.2.2), and empirical results (Section C.2.3).
@@ C.2.3 Empirical Results Comparison
To demonstrate that SDP inherits the same limitations as Diffusion Policy and lacks responsive control, we conduct experiments on the Adroit Relocate task (see Table 16). As the empirical results indicate, Streaming Diffusion Policy performs similarly to Diffusion Policy on the Adroit Relocate task, whereas our method achieves significantly more responsive control than both.
Table 16: We compare Streaming Diffusion Policy with Diffusion Policy and our method on Adroit Relocate task.

### v2

@@ 6.5 Ablation Study
Noise Scheduling Scheme We conduct a comprehensive study to evaluate how different noise scheduling schemes impact the performance of the Responsive Noise-Relaying Diffusion Policy. As shown in Table˜5, the results demonstrate that relying solely on either a linear schedule or a random schedule significantly reduces task success rates. This highlights the importance of integrating mixture scheduling to fully exploit the potential of the noise-relaying buffer design. By combining the two schedules, our model effectively utilizes different noise levels, enhancing the robustness and adaptability of action generation.
Table 5: Ablation study on noise scheduling scheme during training (Section˜5.2). Numbers represent average success rates ($\uparrow$). Numbers in parenthesis indicate the performance drop after removing key component of our Responsive Noise-Relaying Diffusion Policy.
@@
Noise-Relaying Buffer Initialization We conduct an ablation study to evaluate the effectiveness of our initialization scheme. Pure noise directly uses a fully noisy buffer as the initialization for subsequent operations. As shown in Table˜6, laddering initialization achieves significant performance gains compared to pure noise. This result highlights the critical role of our initialization scheme in enabling strong and robust performance.
Table 6: Ablation study on noise-relaying buffer initialization (Section˜5.2).
@@
Noise-Relaying Buffer Capacity We also evaluate the impact of the noise-relaying buffer capacity to the performance, the only tuned hyperparameter in our approach. As shown in Figure˜5, on the Adroit Relocate task, RNR-DP achieves peak performance of $58.5\%$ at a buffer capacity of 84 and maintains strong performance within the range of 56 to 92, demonstrating a wide tolerance. Performance declines when the buffer capacity is too small (e.g., 48) or too large (e.g., 100). Notably, Diffusion Policy achieves only $42.2\%$, while RNR-DP with buffer capacities between 48 and 100 consistently outperforms it. This highlights that for a single task, the noise-relaying buffer offers robust performance over a broad range and is not difficult to tune. Empirically, starting with a value between 56 and 84 and making adjustments suffices most cases.
Model Prediction Type In addition, we investigate the impact of the model’s prediction type. As shown in Table˜7, models predicting the added noise component outperform those directly predicting the action sequence. This finding aligns with the common practice of using noise prediction in diffusion models within the vision domain (Ho et al., 2020).
Table 7: Ablation study on model prediction type (Section˜5.2).
@@ 7 Conclusion and Future Work
Limitations. A limitation of this work is the lack of real-robot evalutions. We discuss our method’s potential advantages in real-robot deployment in Appendix˜E. We will conduct complex real world deployment for future research.
@@ A.1 Task Descriptions
We consider a total of 9 continuous control tasks from 3 benchmarks: ManiSkill2 (Gu et al., 2023), ManiSkill3 (Tao et al., 2024), and Adroit (Rajeswaran et al., 2017). This section provides detailed task descriptions on overall information, task difficulty, object sets, state space, and action space. Some task details are listed in Table˜8.
@@ A.1.1 ManiSkill2 Tasks
  • Overall Description: Pick up a red cube and place it onto a green one. See Figure˜6 for episode visualization.
@@
  • Task Difficulty: This task needs to handle object variations. See Figure˜7 for episode visualization.
@@
  • Overall Description: A dual-arm mobile robot needs to push a swivel chair to a target location on the ground (indicated by a red hemisphere) and prevent it from falling over. The friction and damping parameters for the chair joints are randomized. See Figure˜8 for episode visualization.
@@ A.1.2 ManiSkill3 Tasks
  • Overall Description: It is a simulated version of the real-world push-T task from Diffusion Policy: https://diffusion-policy.cs.columbia.edu/. In this task, the robot needs to precisely push the T-shaped block into the target region, and move the end-effector to the end-zone which terminates the episodes. The success condition is that the T block covers 90% of the 2D goal T’s area. See Figure˜9 for episode visualization.
@@
  • Overall Description: A task where the objective is to push and roll a ball to a goal region at the other end of the table. The success condition is that The ball’s xy position is within goal radius (default 0.1) of the target’s xy position by euclidean distance. See Figure˜10 for episode visualization.
@@ A.1.3 Adroit Tasks
  • Overall Description: The environment is based on the Adroit manipulation platform, a 28 degree of freedom system which consists of a 24 degrees of freedom ShadowHand and a 4 degree of freedom arm. The task to be completed consists on undoing the latch and swing the door open. See Figure˜11 for episode visualization.
@@
  • Overall Description: The environment is based on the Adroit manipulation platform, a 28 degree of freedom system which consists of a 24 degrees of freedom ShadowHand and a 4 degree of freedom arm. The task to be completed consists on repositioning the blue pen to match the orientation of the green target. See Figure˜12 for episode visualization.
@@
  • Overall Description: The environment is based on the Adroit manipulation platform, a 28 degree of freedom ShadowHand and a 4 degree of freedom arm. The task to be completed consists on picking up a hammer with and drive a nail into a board. See Figure˜13 for episode visualization.
@@
  • Overall Description: The environment is based on the Adroit manipulation platform, a 30 degree of freedom system which consists of a 24 degrees of freedom ShadowHand and a 6 degree of freedom arm. The task to be completed consists on moving the blue ball to the green target. See Figure˜14 for episode visualization.
@@ A.2 Demonstrations
This subsection provides the details of demonstrations used in our experiments. See Table˜9. ManiSkill2 and ManiSkill3 demonstrations are provided in Gu et al. (2023) and Tao et al. (2024), and Adroit demonstrations are provided in Rajeswaran et al. (2017).
@@ B.1 Noise-Relaying Diffusion Policy Inference
We summarize the inference pseudo-code of our RNR-DP in Algorithm˜1.
@@
3:  $\mathbf{Q}_{t}$ $\leftarrow\varepsilon_{\theta}(\mathbf{\tilde{Q}}_{t};\mathbf{O}_{t},\{1,\cdots,f\})$ $\triangleright$ $\varepsilon_{\theta}$ is trained using $f$ noise levels
4:  $\mathbf{a}_{t}^{(0)}$ $\leftarrow\mathbf{Q}_{t}\mathrm{.pop}(0)$ $\triangleright$ $\mathbf{a}_{t}^{(0)}$ is a clean action (fully denoised)
5:  $\mathbf{\tilde{Q}}_{t}$ $\leftarrow\mathbf{Q}_{t}\mathrm{.push}(\mathbf{z})$ $\triangleright$ $\mathbf{z}$ is a random noisy action sampled from $\mathcal{N}(\bm{0},\mathbf{I})$
6:  $\mathbf{O}_{t}$ $\leftarrow\mathrm{env.step}(\mathbf{a}_{t}^{(0)})$ $\triangleright$ executes $\mathbf{a}_{t}^{(0)}$ and envrionment updates observation
7:  $\hphantom{1}t$ $\leftarrow t+1$ $\triangleright$ update timestep for the next iteration
@@ B.2 Noise-Relaying Diffusion Policy Training
We summarize the training pseudo-code of our RNR-DP in Algorithm˜2.
@@
6:   $\mathbf{k}=\{k_{1}=1,\cdots,k_{f}=f\}$ $\triangleright$ linear schedule
@@
8:   $\mathbf{k}=\{k_{1}\sim\mathrm{Unif}(\{1,\cdots,f\}),\cdots,k_{f}\sim\mathrm{Unif}(\{1,\cdots,f\})\}$ $\triangleright$ random schedule
@@
10:   $\hat{\mathbf{a}}_{j}=\sqrt{\bar{\alpha}_{t_{j}}}\mathbf{a}_{j}+\sqrt{1-\bar{\alpha}_{t_{j}}}\bm{\epsilon}_{j}$ $\triangleright$ perturbe each $\mathbf{a}_{j}$ independently
@@
[…]
@@ B.3 Policy Architecture
We build our RNR-DP on top of the UNet-based architecture of Diffusion Policy (Chi et al., 2023). The model includes 2 downsampling modules and 2 upsampling modules with each module containing 2 residual blocks. The residual block consists of 1D temporal convolutions (Conv1d), group normalizations (GN), and Mish activation layers. The encoded noise-aware conditioning data (Section˜5.2) is fused into each residual block through the FiLM transformation (Perez et al., 2018). The raw conditioning data is of shape $R^{f\times(C_{\mathrm{emb}}+C_{\mathrm{state}})}$ for state policies and of shape $R^{f\times(C_{\mathrm{emb}}+C_{\mathrm{visual}}+C_{\mathrm{state}})}$ for visual policies. See Figure˜15 for the visualization of a visual policy. We follow the UNet denoiser design for the observation that transformer-based policies are more sensitive to hyperparameters and often require more tuning (Chi et al., 2023). The choice of policy architecture is orthogonal to our method and we believe our design would also improve this policy class.
@@ B.4.1 Key Hyperparameters of RNR-DP
We summarize the key hyperparameters of RNR-DP in Table˜10. The observation horizon $T_{o}$ and noise-relaying buffer capacity $f$ for each task is listed in Table˜11. The number of trainable parameters for each task is listed in Table˜12.
@@ B.4.2 Key Hyperparameters of Diffusion Policy
We summarize the key hyperparameters of Diffusion Policy in Table˜13. The observation horizon $T_{o}$, action executation horizon $T_{a}$ and action prediction horizon $T_{p}$ for each task are listed in Table˜14.
@@ C.1 Empirical Comparison with Acceleration Methods on Visual Observations
We summarize the results of vision-based experiments in Table˜15. As shown in Table˜15, our RNR-DP ourperforms all DDIM variations and CP variations and particularly has an overall improvement over 8-step DDIM by $6.9\%$, over 8-step-chaining CP by $3.4\%$.
Table 15: Evaluation on simpler tasks (Regular Group) not requiring responsive control from ManiSkill and Adroit benchmarks (Visual Observations). We follow our evaluation metric and report values under the same settings as in Table˜4. Tasks in which none of the methods achieve a reasonable success rate under visual observations are omitted.
@@ C.2 Comparsion with Streaming Diffusion Policy (SDP)
Streaming Diffusion Policy (SDP) (Høeg et al., 2024) is a recent advancement over Diffusion Policy that stays close to our approach. In this section, we compare our method with SDP in terms of motivation (Section˜C.2.1), methodology (Section˜C.2.2), and empirical results (Section˜C.2.3).
@@ C.2.3 Empirical Results Comparison
To comprehensively compare Diffusion Policy, Streaming Diffusion Policy, and our method, we conduct experiments on both response-sensitive tasks and regular tasks to evaluate their responsiveness and efficiency, as shown in Table˜16 and Table˜17. The empirical results indicate that on response-sensitive tasks, Streaming Diffusion Policy performs similarly to Diffusion Policy, whereas our method achieves significantly more responsive control than both. On regular tasks, both SDP and our method successfully preserve the performance of DP, but our method achieves 6.25 times faster inference.
Table 16: We compare Diffusion Policy, Streaming Diffusion Policy with our method on response-sensitive tasks.
@@
PushChair w/ g (ManiSkill2) 0.495 0.500 0.547
PushChair w/o g (ManiSkill2) 0.635 0.633 0.694
Table 17: We compare Streaming Diffusion Policy with Diffusion Policy and our method on regular tasks.
StackCube TurnFaucet (w/ g) TurnFaucet (w/o g) Hammer
Method Steps (S) NFEs/a
DP 100 12.5 0.960 0.495 0.595 0.120
SDP 50 6.25 0.961 0.480 0.615 0.136
RNR-DP 1 1 0.935 0.531 0.594 0.139
Appendix D Task Grouping Discussion
In this section, we provide a detailed discussion of the criteria used for grouping tasks into response-sensitive tasks and regular tasks. Section˜D.1 outlines the general rules for task grouping, while Section˜D.2 explains the specific reasoning behind the classification of each task.
D.1 General Rules for Task Grouping
For response-sensitive tasks, inaccurate actions based on outdated observations can easily lead to states not covered by the demonstration dataset. For example, if the robot pushes a swivel chair too forcefully, it may fall and become unrecoverable. In contrast, regular tasks are more tolerant of inaccuracies from outdated observations. For instance, in the stack cube task, even if the gripper doesn’t precisely stop above the cube, this state still falls within the distribution.
D.2 Detailed Task Separation Criteria for Each Task
Relocate (Adroit) (Response-Sensitive) This task requires controlling a high-dimensional dexterous hand to pick up a ball from a surface and transport it to a goal position. Due to the potential for the ball to slip on the surface or shift unpredictably within the fingers, responsive real-time feedback is essential for successful execution.
Door (Adroit) (Response-Sensitive) This task requires controlling a high-dimensional dexterous hand to apply the appropriate force to turn the handle. Insufficient force fails to open the door, while excessive force causes the hand to slip off. Since each phase relies on the precise execution of the previous one, real-time control is essential for success.
PushChair (ManiSkill2) (Response-Sensitive) This task requires controlling a mobile bimanual system to push a swivel chair to a goal position. Effective force control is crucial—pushing too hard can cause the chair to topple over, making recovery impossible. Precisely stopping the chair at the goal position demands highly adaptable control to make fine adjustments. Without real-time feedback, the chair tends to bounce around the goal position instead of coming to an exact stop.
RollBall (ManiSkill3) (Response-Sensitive) This task requires precisely rolling a ball to reach the goal position. Successful execution depends on applying the appropriate force and accurately controlling the rolling trajectory. Even a slight control error can result in missing the target, making highly adaptable and responsive control essential.
PushT (ManiSkill3) (Response-Sensitive) This task involves pushing a T-shaped block to a goal region. Successful completion requires the block to cover 90% of the goal area, necessitating real-time fine adjustments to its position. Without adaptable control, precisely aligning the T-shaped block to the goal position becomes challenging and prone to failure.
StackCube (ManiSkill2) (Regular) This task requires controlling a gripper to pick up a cube and stack it onto another cube. It is tolerant to minor inaccuracies in control, as the gripper does not need to stop precisely above the cube; slight deviations still fall within the expected distribution.
TurnFaucet (ManiSkill2) (Regular) This task requires controlling a gripper to turn on a faucet. Since it involves simple manipulation without complex contact-rich operations, and the faucet remains fixed in place, it is tolerant to inaccuracies in control and does not require real-time adjustments.
Pen (Adroit) (Regular) This task involves controlling a high-dimensional dexterous hand to rotate a pen and align it with a goal pose. Successful completion in the dataset requires an average of only 30 steps, with a minimum of 13 steps, indicating a very short-horizon control requirement. Additionally, numerous studies (Nakamoto et al., 2023; Florence et al., 2022) have identified it as one of the easiest tasks in Adroit.
Hammer (Adroit) (Regular) This task involves controlling a high-dimensional dexterous hand to strike a nail. The hammering motion does not require fine-grained, real-time adjustments, as the primary objective is to deliver sufficient force to the nail. Minor deviations in trajectory or impact position do not significantly impact the task’s success.
Appendix E Complex Real-world Scenarios Discussion
RNR-DP’s responsiveness is crucial in complex real-world scenarios, where dynamic environments, disturbances, and multi-object interactions create significant variability. The ability to quickly adapt to changing conditions is key to achieving robust performance. This responsiveness could enable RNR-DP to tackle real-world challenges like object variations, complex interactions, human disturbances, and environmental uncertainties, broadening its applicability to a wider range of tasks.
We believe RNR-DP offers significant advantages for real-robot deployment. First, it provides responsive control to handle rapid environmental changes, which are crucial in both simulation and real-world scenarios. Second, in real-robot settings, the control policy must meet a specific control rate, unlike in simulation where the environment can wait for policy inference. While DP requires acceleration via DDIM for real-robot deployment, as noted in the DP paper, this does not preserve performance well in our experiments. RNR-DP, on the other hand, better preserves the DP performance while achieving the necessary control rate for real-robot applications.
Appendix F Multi-Modality Property of RNR-DP
To demonstrate the preservation of multi-modality in practice, we visualize action distributions from a state in the ManiSkill2 StackCube task, using our RNR-DP. We sample 1000 actions from our policy then applies PCA dimensionality reduction for visualization. We use histograms to visualize the discrete relative density of these action samples and use kernel density estimation (KDE) to visualize the estimated probability density function. The results are shown in Figure˜16.
Figure 16: Conditional action distribution visualization from our experiments. Our method demonstrates clear bimodal distribution, showing that RNR-DP preserves multi-modality property.
We can see that the dimension-reduced action distribution exhibits a clear bimodal distribution, confirming that multi-modality is preserved in our method.
Appendix G Noise-Relaying Buffer Capacity Discussion
We find that the optimal buffer capacity is closely related to the task horizon, or the number of steps required to complete the task in the dataset. The rationale is that if a task is solved in 30 steps, an 84-step noise-relaying buffer would be inappropriate. Notably, Adroit Pen has the shortest task horizon, an average of 30 steps with a minimum of 13 steps, significantly lower than other tasks, and thus requires a much smaller buffer capacity. In practice, we set the buffer capacity to a value lower than the minimum task horizon in the dataset and make adjustments around this value. For every task, including Adroit Pen, the buffer capacity has a wide range of workable values.
Appendix H Noise Scheduling Discussion
In this section, we aim to discuss two noise scheduling method from Streaming Diffusion Policy (Høeg et al., 2024), 67% Independent 33% Linear and Chunk-Wise Noise Scheduling.
H.1 67% Independent 33% Linear Noise Scheduling
The key difference between Mixture Noise Scheduling and the "67% Independent, 33% Linear" schedule is the proportion of linear noise and random (or independent) noise. In our setup, we use 60% random noise (or independent noise from the SDP paper) and 40% linear noise from TEDi (Zhang et al., 2024), which has proven effective and robust across all tasks. We also tested the "67% independent (random), 33% linear" noise schedule from SDP (Høeg et al., 2024) for the Adroit Relocate task, and the results show it performs well, as the two noise schedules (67-33 vs. 60-40) are quite similar. Refer to Table˜18 for detailed results.
H.2 Chunk-Wise Noise Scheduling
To leverage current observations, RNR-DP uses a single-action rollout during each inference ($T_{a}=1$). Chunk-wise noise scheduling is inappropriate for our setting (Table˜18), as it is designed for action sequence rollouts (e.g., $T_{a}=8$ in DP and SDP). Our scheduling is essentially a special case of chunk-wise noise perturbation, where the chunk size is set to 1.
Table 18: Ablation study on noise scheduling scheme during training. Numbers represent average success rates ($\uparrow$).
Relocate (Adroit)
Ablation
Chunk-Wise 0.001
100% Linear 0.323
100% Random (Independent) 0.389
67% Independent (Random) 33% Linear 0.558
Mixture (60% Random 40% Linear) (Ours) 0.585
Appendix I Mixture Scheduling Visualization
In this section, we provide a detailed discussion of the mechanism and motivation behind Mixture Noise Scheduling. As illustrated in Figure˜17, the random schedule is teaching the model to denoise actions independently, where each action is assigned a random noise level. The linear schedule, on the other hand, maintains an increasing noise level across actions, closely aligning with our inference process through the noise-relaying buffer. Our mixture schedule not only trains the model to denoise actions independently, as in the DP setting, but also ensures smooth transitions between consecutive actions. This better aligns with the noise-relaying buffer structure, resulting in more diverse and robust trainings.
Figure 17: Detailed illustration of the process of Mixture Scheduling.
Appendix J Laddering Initialization Visualization
In this section, we provide a clear and concise discussion of Laddering Initialization. As illustrated in Figure˜18, to transition from random noise to an increased noise level suitable for inference through the noise-relaying buffer, we perform several denoising steps. This process results in a buffer with laddered noise, ensuring a more structured and effective initialization.
Figure 18: Detailed illustration of the process of Laddering Initialization. Everything happens before our policy interacts with the environment.
Appendix K Qualitative Examples of How DP Struggles with Response-Sensitive Tasks and How RNR-DP Resolves Them
Comparison videos of response-sensitive tasks between DP and RNR-DP can be found on this page. The videos clearly demonstrate that while DP struggles with responsive control, RNR-DP effectively completes the tasks with greater precision and adaptability.
Appendix L Discussion on Robosuite Tasks
We initially decided not to experiment with Robosuite for two reasons: (1) nearly all tasks can be solved by DP with close to 100% success, as demonstrated in the DP paper, and (2) the tasks primarily involve pick-and-place, which does not require highly responsive control.
Appendix M Response Sensitivity Analysis with RL Policies
In this section, we assess response-sensitivity through RL training and further validate our task classification. We design two experimental settings. In Action Horizon setting, RL agent directly outputs the next $T_{a}$ actions (as in Diffusion Policy); In Action Repeat setting, RL agent outputs one action which is then executed in the environment repeatedly for $T_{a}$ steps (a strong challenge to responsiveness). We experiment with Adroit Relocate (response-sensitive), Adroit Pen (regular), and ManiSkill2 StackCube (regular) using SAC algorithm.
Figure 19: RL training results on Adroit Relocate (classified as response-sensitive task). Both the Action Horizon and Action Repeat experiments fail to achieve any success on Adroit Relocate, indicating that the task has very low tolerance to outdated actions and strongly depends on the most recent observations. This supports our classification of Adroit Relocate as a response-sensitive task.
Figure 20: RL training results on ManiSkill2 StackCube (classified as regular task). Action Horizon runs can achieve decent success rate, while Action Repeat runs cannot, indicating that ths task has some tolerance to outdated actions. This supports our classification of ManiSkill2 StackCube as a regular task.
Figure 21: RL training results on Adroit Pen (classified as regular task). Both Action Horizon and Action Repeat runs have reasonable success rate, indicating that the task has very high tolerance to outdated actions. This supports our classification of Adroit Pen as a regular task.
In the Action Horizon setting, $T_{a}=8$ is unsolvable due to the large action space, which hinders RL exploration. Behavior Cloning handles this better since it’s supervised. Hence, we focus on $T_{a}=4$, where Adroit Pen and ManiSkill2 StackCube achieve reasonable success, while Adroit Relocate struggles and achieves nearly-zero success rate (See Figures˜19, 21 and 20). In the Action Repeat setting, repeating the same action for $T_{a}=8$ steps poses a high responsiveness challenge. Even so, Adroit Pen still achieves some success, showing strong tolerance to outdated actions (See Figure˜21). In summary, these experiments confirm that Adroit Pen and ManiSkill2 StackCube tolerate outdated actions better than Adroit Relocate, supporting our task classification.